What is the maximum number of electrons in an atom that can have the following quantum numbers?
(a) n = 2, ms = -1/2



(b) n = 6, l = 3



(c) n = 7, l = 5, ml = +3



(d) n = 4, l = 2, ml = +2, ms = -1/2
1
Expert's answer
2018-04-05T07:38:11-0400
(a) The total electrons in n is given by 2 n2, n = 2, number of electrons = 2 * 2^2 = 8
half of 8 electrons will have spin quantum number ms = -1/2 i.e. 4 electrons
(b) 14 electrons
(c) 2 electrons
(d) 1 electron